The lenght of a rectangle is one inch less than twice its width. The diagonal of the rectangle is two inches more than its length. Find the area of the rectangle.
Answer by edjones(8007)

L=2W-1
c=diagonal and is the hypotenuse of a triangle with sides L and W
c=2W+1
a^2+b^2=c^2
(2W-1)^2+W^2=(2W+1)^2
4W^2-4W+1+W^2=4W^2+4W+1
W^2-8W=0
W(W-8)=0
W=0, W=8 Since the rectangle can't have a width of 0 W=0 is not an answer.
W=8 is the only answer.
L=15
c=17
A=15*8=120sq inches (Answer)
check:
8^2+15^2=64+225=289
17^2=289
Ed